OpenStudy (tal_tal):

What is the value of the expression shown below? 3 over 5 to the power of 2 + 4 × 3 − 2 10 and 9 over 25 11 and 2 over 25 12 and 9 over 25 13 and 2 over 25

rebeccaxhawaii (rebeccaxhawaii):

2 + 12 - 2 = 12

rebeccaxhawaii (rebeccaxhawaii):



OpenStudy (anonymous):

or maybe \[\left(\frac{3}{5}\right)^2+stuff\]

OpenStudy (mathstudent55):

It's very hard to know what you mean. Is this it?